OpenStudy (anonymous):

Find y': y = (2x - 2)(3x + 1)

OpenStudy (mathteacher1729):

It's probably easier to just expand the binomial and then differentiate term-by term. \[y = 6x^2-4x-1\] Differentiating that should be fairly easy.

OpenStudy (anonymous):

are you please able to show the steps? i get that u expand the brackets and so on...

OpenStudy (anonymous):

You need help with expanding from the parentheses or finding the derivative from y = 6x^2 - 4x - 1?

OpenStudy (anonymous):

expanding the parentheses

OpenStudy (anonymous):

(2x - 2)(3x +1), you use the distributive property. The 3x + 1 is distributed to the 2x and to the -2... so (2x - 2)(3x+1) = 2x(3x+1) - 2(3x + 1) = 6x^2 + 2x - 6x + 1 = 6x^2 - 4x - 2... take the derivative of 6x^2 - 4x - 2.

OpenStudy (anonymous):

great thank you heaps!!!

OpenStudy (radar):

Wouldn't the actual expansion be:\[y=6x ^{2}-4x-2\]

OpenStudy (radar):

Oh I see where gyasbic pointed that out, it wouldn't affect the answer anyway....would it

OpenStudy (anonymous):

the final answer is 12x - 4